# Classical Simulations of Low Magic Quantum Dynamics

### v1

@@ B.1 Cluster
Figure 11: (a) The first moment of the maximal cluster size $\overline{n_{\mathrm{max}}}$ as a function of the measurement probability $p_{m}$. $\overline{\phantom{X}}$ denotes averaging over circuits. Data is shown for the single-pair all-to-all model with $X$-basis measurements. (b) Rescaled first moment of the maximal cluster size $\overline{n_{\mathrm{max}}}L^{-\beta/\nu}$ as a function of the measurement probability $p_{m}$, which exhibits a crossing at the critical connectivity (cc) rate $p_{m}=p_{m}^{cc}$ (Inset) Scaling collapse for maximal cluster size used to obtain the critical exponents in Table 1. (c) Normalized variance of the peak cluster size, $\overline{n_{\mathrm{max}}^{2}}/\overline{n_{\mathrm{max}}}^{2}$ as a function of measurement probability $p_{m}$. The location of the peak is a proxy of $p_{m}^{cc}$.
@@ B.2 Entanglement Transition
Figure 13: (a) Scaling of $\overline{S}_{mm,1}$ with $L$ for $p_{m}\in\{0.13,0.2,0.27,0.34,0.41,0.48,0.55,0.62\}$. We find the critical entanglement (ce) rate $p_{m}^{ce}=0.27(1)$ as indicated by the inflection point. (b) Max-min entanglement entropy $\overline{S}_{mm,1}$ as a function of measurement probability $p_{m}$. $\overline{\phantom{X}}$ denotes averaging over circuits. Data is shown for system sizes $L\in\{8,12,14,16,18,20\}$ in the single pair all-to-all model with $X$-basis measurements. The crossing suggests the location of the transition (Inset) $\overline{S}_{mm,1}$ as a function of $L$, which exhibits log scaling at the transition.
@@ B.3.1 Fitting Procedure
Figure 14: Late time decay of $\overline{S_{\text{Q}}}$, which is the entanglement entropy between the ancilla qubit and the system. $\overline{\phantom{X}}$ denotes averaging over circuits. This shows the exponential decay regime used to extract the decay time $\tau$, where $\overline{S_{\text{Q}}}\sim e^{-t/\tau}$. Data is shown for the single-pair all-to-all model with $X$-basis only measurements at the critical purification rate $p_{m}^{cp}=0.26(1)$ for the purification transition.
@@ B.3.2 Additional Result on Purification Transition
We now present data supporting the phase boundary of the purification transition (5) in the single-pair all-to-all model with $X$-basis only measurements. Figures 15 and 16 show the scaling behavior of the decay time $\tau$ as a function of both system size $L$ and measurement probability $p_{m}$ for different $T$-gate densities. These results provide further confirmation that the purification transition persists even in the presence of non-Clifford gates, and allow for an independent extraction of the dynamical exponent $z$ and the critical purification rate $p_{m}^{cp}$.
Figure 15: Dependence of the decay time $\overline{\tau}$ as a function of system size $L$, where $\overline{S_{Q}}\sim e^{-t/\tau}$ and $S_{Q}$ is the entanglement between the ancilla and the system. $\overline{\phantom{X}}$ denotes averaging over circuits. Data is shown for the single-pair all-to-all model with $X$-basis only measurements. The measurement probabilities $p_{m}$ are obtained near the critical purification (cp) rate $p_{m}^{cp}$. We extract $p_{m}^{cp}$ and $z$, where $\tau(L,p_{m}^{cp})\sim L^{z}$, by looking for the inflection point in this family of curves, and fitting the slope. The probability of a $T$-gates is given as (a) $p_{T}=0$, (b) $p_{T}=2/L^{2}$, (c) $p_{T}=4/L^{2}$, and (d) $p_{T}=8/L^{2}$. The fit used to obtain z is shown as a dashed red line.
Figure 16: Dependence of the decay time $\overline{\tau}$ as a function of the measurement probability $p_{m}$, where $\overline{S_{Q}}\sim e^{-t/\tau}$ and $S_{Q}$ is the entanglement between the ancilla and the system. $\overline{\phantom{X}}$ denotes averaging over circuits. Data is taken in the single-pair all-to-all model with $X$-basis only measurements. Here $z\approx 1/5$, where $\tau(L,p_{m}^{cp})\sim L^{z}$, is the dynamical critical exponent fixed from the inflection point analysis at the critical purification (cp) rate, consistent with the crossing. The probability of a $T$-gates is given as (a) $p_{T}=0/L^{2}$, (b) $p_{T}=2/L^{2}$, (c) $p_{T}=4/L^{2}$, and (d) $p_{T}=8/L^{2}$.

### v2

@@ B.1 Cluster
Figure 11: (a) The first moment of the maximal cluster size $\overline{n_{\mathrm{max}}}$ as a function of the measurement probability $p_{m}$. $\overline{\cdots}$ denotes averaging over circuit trajectories. Data is shown for the single-pair all-to-all model with $X$-basis measurements. (b) Rescaled first moment of the maximal cluster size $\overline{n_{\mathrm{max}}}L^{-\beta/\nu}$ as a function of the measurement probability $p_{m}$, which exhibits a crossing at the critical connectivity (cc) rate $p_{m}=p_{m}^{cc}$ (Inset) Scaling collapse for maximal cluster size used to obtain the critical exponents in Table 1. (c) Normalized variance of the peak cluster size, $\overline{n_{\mathrm{max}}^{2}}/\overline{n_{\mathrm{max}}}^{2}$ as a function of measurement probability $p_{m}$. The location of the peak is a proxy of $p_{m}^{cc}$.
@@ B.2 Entanglement Transition
Figure 13: (a) Scaling of $\overline{S}_{mm,1}$ with $L$ for $p_{m}\in\{0.13,0.2,0.27,0.34,0.41,0.48,0.55,0.62\}$. We find the critical entanglement (ce) rate $p_{m}^{ce}=0.27(1)$ as indicated by the inflection point. (b) Max-min entanglement entropy $\overline{S}_{mm,1}$ as a function of measurement probability $p_{m}$. $\overline{\cdots}$ denotes averaging over circuit trajectories. Data is shown for system sizes $L\in\{8,12,14,16,18,20\}$ in the single pair all-to-all model with $X$-basis measurements. The crossing suggests the location of the transition (Inset) $\overline{S}_{mm,1}$ as a function of $L$, which exhibits log scaling at the transition.
@@ B.3.1 Fitting Procedure
Figure 14: Late time decay of $\overline{S_{\text{Q}}}$, which is the entanglement entropy between the ancilla qubit and the system. $\overline{\cdots}$ denotes averaging over circuit trajectories. This shows the exponential decay regime used to extract the decay time $\tau$, where $\overline{S_{\text{Q}}}\sim e^{-t/\tau}$. Data is shown for the single-pair all-to-all model with $X$-basis only measurements at the critical purification rate $p_{m}^{cp}=0.26(1)$ for the purification transition.
@@ B.3.2 Additional Result on Purification Transition
We now present data supporting the phase boundary of the purification transition (Fig. 5) in the single-pair all-to-all model with $X$-basis only measurements. Figures 15 and 16 show the scaling behavior of the decay time $\tau$ as a function of both system size $L$ and measurement probability $p_{m}$ for different $T$-gate densities. These results provide further confirmation that the purification transition persists even in the presence of non-Clifford gates, and allow for an independent extraction of the dynamical exponent $z$ and the critical purification rate $p_{m}^{cp}$.
Figure 15: Dependence of the decay time $\overline{\tau}$ as a function of system size $L$, where $\overline{S_{Q}}\sim e^{-t/\tau}$ and $S_{Q}$ is the entanglement between the ancilla and the system. $\overline{\cdots}$ denotes averaging over circuit trajectories. Data is shown for the single-pair all-to-all model with $X$-basis only measurements. The measurement probabilities $p_{m}$ are obtained near the critical purification (cp) rate $p_{m}^{cp}$. We extract $p_{m}^{cp}$ and $z$, where $\tau(L,p_{m}^{cp})\sim L^{z}$, by looking for the inflection point in this family of curves, and fitting the slope. The probability of a $T$-gates is given as (a) $p_{T}=0$, (b) $p_{T}=2/L^{2}$, (c) $p_{T}=4/L^{2}$, and (d) $p_{T}=8/L^{2}$. The fit used to obtain z is shown as a dashed red line.
Figure 16: Dependence of the decay time $\overline{\tau}$ as a function of the measurement probability $p_{m}$, where $\overline{S_{Q}}\sim e^{-t/\tau}$ and $S_{Q}$ is the entanglement between the ancilla and the system. $\overline{\cdots}$ denotes averaging over circuit trajectories. Data is taken in the single-pair all-to-all model with $X$-basis only measurements. Here $z\approx 1/5$, where $\tau(L,p_{m}^{cp})\sim L^{z}$, is the dynamical critical exponent fixed from the inflection point analysis at the critical purification (cp) rate, consistent with the crossing. The probability of a $T$-gates is given as (a) $p_{T}=0/L^{2}$, (b) $p_{T}=2/L^{2}$, (c) $p_{T}=4/L^{2}$, and (d) $p_{T}=8/L^{2}$.